if $2000 is invested at a rate of 5% compound continuously, approximate the balance in the account after 20 years. Use the formula A=Pe^ rt
Answer by rfer(16322)

A=2000(2.7183)^0.05*20
A=2000(2.7183)^1
A=2000*2.7183
A=$5436.60